### v1

# MiM: Mask in Mask Self-Supervised Pre-Training for 3D Medical Image Analysis

Abstract
The Vision Transformer (ViT) has demonstrated remarkable performance in Self-Supervised Learning (SSL) for 3D medical image analysis. Mask AutoEncoder (MAE) for feature pre-training can further unleash the potential of ViT on various medical vision tasks. However, due to large spatial sizes with much higher dimensions of 3D medical images, the lack of hierarchical design for MAE may hinder the performance of downstream tasks. In this paper, we propose a novel Mask in Mask (MiM) pre-training framework for 3D medical images, which aims to advance MAE by learning discriminative representation from hierarchical visual tokens across varying scales. We introduce multiple levels of granularity for masked inputs from the volume, which are then reconstructed simultaneously ranging at both fine and coarse levels. Additionally, a cross-level alignment mechanism is applied to adjacent level volumes to enforce anatomical similarity hierarchically. Furthermore, we adopt a hybrid backbone to enhance the hierarchical representation learning efficiently during the pre-training. MiM was pre-trained on a large scale of available 3D volumetric images, i.e., Computed Tomography (CT) images containing various body parts. Extensive experiments on thirteen public datasets demonstrate the superiority of MiM over other SSL methods in organ/lesion/tumor segmentation and disease classification. We further scale up the MiM to large pre-training datasets with more than 10k volumes, showing that large-scale pre-training can further enhance the performance of downstream tasks. The improvement also concluded that the research community should pay more attention to the scale of the pre-training dataset towards the healthcare foundation model for 3D medical images.
keywords:
[…]
journal: Preprint
1 Introduction
Fig. 1: Different SSL for 3D medical image analysis. Current Masked Image Modeling methods for 3D medical images primarily (a) rely on pretext tasks e.g., inpainting, at a single level, utilizing hybrid transformers to incorporate all tokens or (b) employ an MAE that reconstructs at a single level using visible tokens. In contrast, (c) we observe that 3D medical images inherently exhibit hierarchical properties. Thus, our Mask in Mask (MiM) framework aims to encode multi-level 3D medical image learning across hierarchical visual tokens at various scales through multi-level reconstruction and cross-level alignment (we set the number of level $L$ to 3 in this figure). Additionally, our framework employs a hybrid transformer while only using visible tokens.
Deep learning has made significant progress in medical image analysis, albeit limited to the supervised learning paradigm that relies on a relatively large amount of labeled training data [57, 93]. However, labeling 3D medical images is a time-consuming and labor-intensive process that requires expertise [39, 60]. Specifically, to expand the training set while reducing human intervention, Semi-supervised Learning and Self-Supervised Learning (SSL) has been become a promising direction  [16, 30, 31, 10, 74]. Among them, SSL has shown great potential in learning transferable representations from unlabeled data without any annotations, which can be further fine-tuned on downstream tasks to enhance the task performance [38, 34, 62, 78, 70, 91, 89, 94].
MAE-based methods [30, 6, 64, 21, 66, 67, 46, 79] has achieved high performance in various natural image analysis tasks by learning transferable representations through reconstructing the original output in pixels from its highly-masked version based on the ViT structure. However, 3D medical images such as Computed Tomography (CT), which are often volumetric and in a much higher dimension, contain a multi-granularity of visual details. Existing methods [38, 34, 62, 78, 70, 91, 89, 94, 76, 69, 72] like Swin-UNETR[62] proposes to learn local and global representation by combining tailored proxy tasks including inpainting, rotation, and contrastive different views. He et al. [34] further explores the geometric similarity between global and local crops by affirmation and matching. Jiang et al. [38], on the other hand, maximizes the similarity between different views that contain consistent semantics. MAE3D [16] reconstructs the missing pixel based on a cropped volume. These methods mainly focus on encoding the 3D medical image from local or global perspectives at a single scale, which may still result in the loss of finer granularity detail.
In the case of medical images in large spatial sizes of high dimensions, the use of hierarchical design is crucial for effective analysis and interpretation of clinical information [9, 62, 34, 35]. In this paper, we propose a novel framework that enhances representation learning of MAE for 3D medical images by hierarchical design, as illustrated in Fig. 1. To achieve this goal, our framework utilizes a hierarchical view of 3D medical images, containing rich multi-granularity of clinical information essential for further analysis and providing important clinical insights. To obtain hierarchical masked volumetric inputs, we propose a multi-level volume generation strategy. This strategy conveys information from a coarse to a finer level, which combines to bring forth a rich representation of the model. We employ multi-level reconstruction to simultaneously reconstruct the masked volumes with multi-granularity details to learn discriminative representation. Compared to single-level reconstruction [16, 66, 79, 22], multi-level reconstruction can capture more detailed information and improve representation learning [67]. Our multi-levels are constructed by dividing the original volume into multiple levels, which is different from the multi-scale strategy that uses different scales of the same volume [67, 35]. Cross-level alignment between adjacent level volumes is applied to learn both the global context information and local details, achieving better representation. Positive pairs i.e., volumes from the same original volumes but at the adjacent level granularity, share the same clinical information. We enable the model to learn more robust representations by aligning the features of these volumes at different scales. We extend the hybrid backbone for MAE [22] from 2D natural images to our 3D medical image pre-training framework to further improve the efficiency during the pre-training period. Since previous MAE-based works [79] require all tokens to be kept in the later transformer stages, leading to an unaffordable computation cost. This leads to the loss of MAE’s efficiency advantage.
In summary, our contribution can be categorized into three aspects:
  1. 1.
    We propose a simple yet effective SSL framework, called MiM, which advances MAE with hierarchical design in 3D medical image pre-training, pushing the representability of pre-trained models.
  2. 2.
    MiM encodes multi-level visual cues across various scales into the metric, providing a reliable measurement of the hierarchical structure of 3D medical images. A multi-level volume generation strategy is proposed, and two tailored proxy tasks, multi-level reconstruction, and cross-level alignment, are presented, which simultaneously learn powerful local and global representations.
  3. 3.
    We pre-train MiM with various sizes of pre-training datasets ranging from 1k to 10k volumes. Extensive experiments demonstrate its effectiveness and achieve state-of-the-art performance across various tasks on thirteen public datasets. We conclude that scaling up the pre-training dataset can further improve the performance of MiM.
2 Related works
In this section, we first review the previous mainstream self-supervised learning paradigms, especially for existing SSL methods for 3D medical images. Then we introduce the recent advances in masked image modeling and hierarchical design of SSL for 3D medical images for comparison with our method and highlight the differences.
SSL for Medical Image Analysis. Previous studies in SSL can be categorized into two primary paradigms: Contrastive learning and Pretext task [31, 10, 23]. Contrastive learning methods bring positive pairs of images (e.g., different views of the same image) closer and separate negative pairs (e.g., different images) away in the feature space [11, 10, 31, 12, 15]. For example, MoCo v3 utilized a memory bank to maintain consistent representations of negative samples during contrastive learning [31, 13, 15]. These works have also shown effectiveness in medical image analysis, such as dermatology classification and chest X-ray classification [87, 90, 80]. However, previous contrastive learning approaches have primarily focused on the semantic misalignment of different instances, leading to minor improvements in dense prediction, such as segmentation [5, 75, 84, 90]. Pretext task, on the other hand, are designed to explore the inner space structure information of images and have shown promise in dense prediction tasks [7, 26, 63, 77]. For example, Zhou et al. [91] proposed to directly pre-train a CNN model by restoring volumetric medical images from their distorted versions. Zhu et al. [92] utilized the concept of solving a Rubik’s Cube to learn structural features from the volumetric medical data. Additionally, five common pretext tasks have been verified to be effective for volumetric image pre-training [61]. Tang et al. [62] proposed to pre-train a 3D swin transformer [48] with the combination of three pretext tasks. Recently, Masked Image Modeling [30, 79], considered as widely used pretext tasks based, shows its effectiveness on 3D medical analysis [16, 76, 94].
Masked Image Modeling. Masked Image Modeling is a generative SSL technique that involves training on images that have been corrupted by masking [30, 79, 16, 6, 66, 70, 76, 69, 94]. Previous researches on this technique, such as Vincent et al. [65] and Pathak et al. [52], treated masking as a type of noise and used inpainting techniques to predict the missing pixels. Dosovitskiy et al. [18] investigated masked patch prediction by predicting the mean color of images. He et al. [30] adhered to the spirit of raw pixel restoration and demonstrated that masking a high proportion of input images can yield a non-trivial and meaningful self-supervisory task.  Chen et al. [6], Wang et al. [66] proposes to improve masking priority strategies [66] and the reconstruction targets [6]. Zhang et al. [81], Wang et al. [69] introduce the concept of Masked Image Modeling to versetile 2D medical scenario e.g. disease classfication [76] and segmentation [69]. Chen et al. [16] represented a recent advancement in Masked Image Modeling [30, 79] for 3D CT in medical image analysis. Zhuang et al. [94] proposed to advance MAE for 3D medical image analysis, from the global and local perspectives by incorporating the contrastive learning paradigm.
Our method follows the primary idea of Masked Image modeling. As shown in Fig. 1, to address the existing problems mentioned above, instead of single-level masking, MiM leverages the hierarchical design inherent in the 3D medical images to generate multi-level volume for reconstruction and cross-level alignment, which can effectively capture the hierarchical structure of 3D medical images and improve the representation learning.
Hierarchical SSL. Hierarchical SSL has been investigated in numerous studies [79, 6, 22, 83] in the context of natural images.  Xie et al. [79] and  Chen et al. [6] proposed to incorporat Masked Image Modeing [30] with a hierarchical backbone e.g., Swin Transformer [48]. Wang et al. [67] reconstructing masked patches on multi-scale features using ViT [19], while  Gao et al. [22] advocated for predicting masked patches using only visible patches and a hierarchical backbone. Hosseinzadeh Taher et al. [35] divided single-scale images into multi granularities and aligned their contrastive views [31, 12] with ResNet [32].
However, the relatively low dimensionality of 2D images limits the effectiveness of learning the hierarchical structure of 3D medical images given a single-level image with hybrid features. Furthermore, previous methods [35, 67] treated each level image separately, neglecting the cross-level shared semantics.
In this paper, we propose a novel pretext task of Masked Image Modeling with a hierarchical design tailored for 3D medical images. This approach considers the inherent hierarchical design of 3D medical images during pre-training. Distinct from previous methods, we introduced a totally different Masked Image Modeling paradigm for 3D medical images. Our proposed hierarchical Masked Image Modeling paradigm employs multi-level volume generation for reconstruction and cross-level alignment, effectively capturing the hierarchical structure of 3D medical images and enhancing representation learning.
3 Methodology
This section provides an overview of our proposed MiM method. Firstly, the overall framework of the MiM method is introduced in Section 3.1. Secondly, the process of multi-level reconstruction is presented in Section 3. Then, the cross-level alignment via contrastive learning in our proposed MiM method is described in Section 3.3. Finally, the backbone of the MiM method in the pre-training period is introduced in Section 3.4.
Fig. 2: The overall framework of our MiM pre-training framework. The level $L$ is set to 3 for better illustration. We first conduct the process of multi-level masked volume generation. The multi-level reconstruction module is responsible for reconstructing the masked volumes at different levels. The cross-level alignment module aligns representations of volumes between volumes from adjacent levels, aiming to enforce anatomical similarity in a hierarchical manner.
3.1 Overall framework
The proposed MiM method’s overall framework is presented in Fig. 2, which comprises multi-level reconstruction modules and cross-level alignment modules. To pre-train the model with MiM, we first generate multi-level volumes from the input 3D medical images. Then, an input volume is cropped into non-overlapping patches, which are divided into unmasked patches and masked patches. The unmasked patches are transformed into a high-dimension feature space using a typical backbone (CNN [32] and transformer [18]). The masked patches are used to generate the next level of masked volumes for multi-level learning. The goal is to restore the masked patches from different levels of masked volumes. In this paper, instead of a single level of masked volume [21, 66, 67], we propose to ease this goal by multi-level of masked volume. We developed a $\mathcal{L}_{R}$ to supervise the final prediction. In addition, we further use a loss function $\mathcal{L}_{C}$ to align the shared semantics between the cross-level volumes, aiming to learn the global semantic content as well as local details. Further details are presented in Section 3 and Section 3.3.
3.2 Multi-level reconstruction
Generation of multi-level masked volumes. Given a volumetric image $x\in\mathbb{R}^{CHWD}$ (e.g., $C=1$ for CT), we aim to generate multi-level volumes $\{x^{l}\in\mathbb{R}^{CH^{l}W^{l}D^{l}},l\in L\}$, which contains $L$ numbers of level, i.e., different granularity information from coarse to fine level. The generation of multi-level volume is illustrated in Fig. 3. We start by cropping a large portion of sub-volume from $x$ as Level-1 volume $x^{1}\in\mathbb{R}^{CH^{1}\times W^{1}\times D^{1}}$. The level 1 volume $x^{1}$ is then patchified into $N$ non-overlapping visual tokens using a patch of $\{\frac{H^{1}}{6},\frac{W^{1}}{6},\frac{D^{1}}{6}\}$ as in  [30, 16, 94]. Following previous MAE-based methods [16, 94], we apply a high mask ratio $\mu$ (e.g., 60%) to the $N$ non-overlapping visual tokens (e.g., 216), resulting in visible tokens and invisible tokens. Each visible token is resized at spatial dimension to $\mathbb{R}^{Chwd}$ before further processed by a linear project as previous MAE-based methods [16, 94]. The invisible tokens from Level-1 $x^{1}$ are considered as the next level volume, i.e., Level-2 volume $x^{2}\in\mathbb{R}^{CH^{2}\times W^{2}\times D^{2}}$. It’s noted that the Level-2 volume $x^{2}$ is generated from the masked patches of Level-1 volume $x^{1}$, instead of unmasked patches. The content of the target reconstruction at different levels shares regions, meaning that the volumes for Level-1 and Level-2 reconstruct overlapping regions but from different granularities, respectively. This can effectively capture the hierarchical structure of 3D medical images and improve representation learning [67]. Ablation study of the reconstruction targets, i.e., the choices of masked patches or unmasked patches as next level volume, is presented in Section 4.7.1. Since there exist many invisible tokens, we randomly sample $\gamma$ times without replacement from these tokens for computation efficiency.
Fig. 3: Illustration of the multi-level masked volume generation. The level $L$ is set to 2 and slices drawn from the 3D medical images are shown in the figure for better illustration. Level-2 volume $x^{2}$ is randomly sampled from the masked patches of Level-1 volume $x^{1}$ (i.e., patch with red box). Consequently, the content of the target reconstruction at different levels shares regions, meaning that the volumes for Level-1 and Level-2 reconstruct overlapping regions but from different granularities, respectively.
The Level-2 volume $x^{2}$ is also patchified into $N$ non-overlapping visual tokens with a patch size of $\{\frac{H_{2}}{6},\frac{W_{2}}{6},\frac{D_{2}}{6}\}$. With applying the mask ratio $\mu$ as in the previous process, the visible tokens from Level-2 volume $x^{2}$ are resized to $\mathbb{R}^{Chwd}$ before feeding the backbone. Meanwhile, the invisible tokens from Level-2 volume $x^{2}$ are considered as next finer level volume i.e., Level-3 volume $x^{3}\in\mathbb{R}^{CH^{3}\times W^{3}\times D^{3}}$. Similarly, we randomly sample $\beta$ times without replacement from these tokens for computation efficiency. This generation process of multi-level volume is repeated until the last level of volume $x^{L}$ is generated.
Thus, all visible tokens from different levels are fed into the backbone to extract the high-dimension features $z$. Following previous SSL methods [30, 16, 94], a lightweight decoder is used to project $z$ along with learnable invisible tokens into latent features $q$. For reconstruction, the decoder followed by a simple prediction head is used to reconstruct the invisible tokens $y$. For cross-level alignment, all visual tokens from identical levels are aggregated by MLP layers and then projected into a context vector $c$.
Single-level reconstruction.  Our reconstruction target is the pixel values of the invisible tokens for each level volume. With feature extracted from the backbone and the decoder with a predition head (i.e., a linear layer), following previous MAE methods [16, 94], we reshape the prediction results and reconstruction targets to one-dimensional vector, i.e., $\hat{y}\in\mathbb{R}^{{1\times C}}$ and $y\in\mathbb{R}^{{1\times C}}$, where $C$ is the number of dimensions. Specifically, we empirically set the $C$ to 768, as in  [30, 16, 94].
Then we compute the differences $d$ between the reconstruction target $y$ and the prediction result $\hat{y}$. Specifically, we use MSE distance [30, 16] to measure the difference $d_{m}$ between the prediction result and the reconstruction target for each masked token $m$ as follows:
$d_{m}=\|y_{m}-\hat{y}_{m}\|_{2},m\in M,M=\mu N,$ (1)
where $M$ denotes the numbers of masked tokens and $d_{m}$ ranges from 0 to 1. Lower $d_{m}$ indicates the smaller difference between the prediction result $\hat{y}_{m}$ and the reconstruction target $y_{m}$.
Aiming to minimize the difference $d_{m}$, we define the reconstruction loss $\mathcal{L}_{\mathcal{R}}^{l}$ for each level of masked volumetric images $l\in L$ as follows:
$\mathcal{L}_{\mathcal{R}}^{l}=\frac{1}{|M|}\sum_{m=1}^{|M|}d_{m}^{l},l\in L,$ (2)
where $M$ represents the number of mask tokens. The reconstruction loss is computed as the sum of squared differences between the reconstruction target and the reconstructed representations based on pixel values.
Loss function for multi-level reconstruction. To learn multi-granularity details from 3D volumetric images, We apply multi-level reconstruction for each level of masked volumetric images, ranging from coarser levels to finer levels. The formulation of the multi-level reconstruction loss $\mathcal{L}_{\mathcal{R}}$ is shown as follows:
$\mathcal{L}_{\mathcal{R}}=\frac{1}{L}\sum_{l=1}^{L}\mathcal{L}_{\mathcal{R}}^{l}$ (3)
3.3 Cross-level alignment
The alignment between the shared semantic patches of cross-level volumes from fine-to-coarse can enforce anatomical similarity in a hierarchical manner. Fig. 4 illustrates the process of cross-level alignment. Since we generate the finer level volumes from the coarser level volumes (e.g., Level-2 volumes generated from masked patches of the Level-1 volumes), these volumes must share semantic context, which can be regarded as positive pairs. In contrast, the non-overlap patches in the coarser volumes (e.g., the rest patches in the Level-1 volumes) are considered as negative patches. To enlarge the high-dimensional feature consistency for shared anatomical structure patches (i.e., positive pairs) and discrepancy between non-overlap patches (i.e., negative pairs), we apply contrastive learning [10] to the context vector $c$ and patches $p$ Specially, we reshape the context vector and patches its coarser volumes to one-dimension vector $c\in\mathbb{R}^{1\times D}$ and $\{p_{i}\in\mathbb{R}^{1\times D},i\in N\}$, where $D$ is the number of dimensions, empirically set $D$ to 2048, as in  [4].
Fig. 4: Illustration of cross-level alignment module. We set level $L$ to 2 for better illustration. The Level-2 context representation and the corresponding patches of Level-1 volumes are regarded as positive pairs. The rest patches of the Level-1 volume are considered negative pairs.
First, given the context vector $c_{i}$ and the patches $p$, we compute the cosine similarity $s_{ij}$ between the context vector $c_{i}$ and the patches $p_{j}$ as follows:
$s_{ij}=cosSim(c_{i},p_{j})=\frac{c_{i}\cdot p_{j}}{\|c_{i}\|\ \|p_{j}\|},i\in M,j\in N.$ (4)
We aim to maximize the cosine similarity between the context vector $c_{i}$ and the positive patches $p$ and minimize the cosine similarity between the context vector $c_{i}$ and the negative patches $p_{j}$ from the rest patches of the volumes. Thus, we apply the contrastive loss [14] to implement this goal as follows,
$\ell_{ij}=-\textrm{log}\frac{\textrm{exp}\ (s_{ij}/\tau)}{\sum_{k=1}^{N}\mathbbm{1}_{k\neq i}\ \textrm{exp}(s_{ik}/\tau)},i\in M,j\in N,M\in\mu N.$ (5)
Thus, the cross-level alignment loss $\mathcal{L}_{\mathcal{C}}^{l,l+1}$ between adjacent level $l$ and $l+1$ volumes is computed as follows:
$\mathcal{L}_{\mathcal{C}}^{l,l+1}=\frac{1}{M\cdot N}\sum_{i=1}^{M}\sum_{j=1}^{N}l_{ij},l\in L-1,M\in\mu N.$ (6)
The cross-level alignment loss is computed as the sum of the negative log-likelihood of the cosine similarity between the context vector and the rest patches from the coarser volumes.
The formulation of cross-level alignment $\mathcal{L}_{\mathcal{C}}$ is defined as the average of the cross-level alignment loss between adjacent level volumes, as shown in Eq. 7,
$\mathcal{L}_{\mathcal{C}}=\frac{1}{L-1}\sum_{l=1}^{L-1}\mathcal{L}_{\mathcal{C}}^{l,l+1}$ (7)
By minimizing the overall cross-level alignment loss, we can enforce anatomical similarity hierarchically.
3.4 Backbone
Previous hybrid transformers, such as nnFormer [88] and Swin-UNETR [28, 62], aim to produce pyramid features. However, they suffer from complexity, needing to process all visual tokens for Masked Image Modeling, which is inefficient [22] and hinders scalability. We extend the backbone of [22] from 2D to 3D medical images by adding an extra dimension channel i.e., depth. This modification allows us to process visual tokens only on transformer layers, enhancing computational efficiency and scalability. The ablation study confirming the effectiveness of this approach is presented in Section 4.7.1.
Overall objective function. Our proposed MiM framework, as illustrated in Algorithm 1, introduces a hierarchically designed approach to 3D medical image representation learning through multi-level reconstruction $\mathcal{L}_{R}$ and cross-level alignment $\mathcal{L}_{C}$. We empirically set the level of masked volume to $L=3$ in this paper since the three-level masked volume can provide a good balance between the representation learning and computational efficiency. Abaltion study of the level of masked volume $L$ is on the Section 4.7.1. Specifically, the multi-level reconstruction loss in Eq. 2 can be further expanded as follows:
$\mathcal{L}_{\mathcal{R}}=\mathcal{L}_{\mathcal{R}}^{1}+\mathcal{L}_{\mathcal{R}}^{2}+\mathcal{L}_{\mathcal{R}}^{3}.$ (8)
Then, since cross-level alignment loss applied between adjacent level volumes, Eq. 6 can be also further expanded as follows:
$\mathcal{L}_{\mathcal{C}}=\mathcal{L}_{\mathcal{C}}^{1,2}+\mathcal{L}_{\mathcal{C}}^{2,3}.$ (9)
Thus, the total loss function $\mathcal{L}$ is the combination of these two losses, as shown in Eq. 10,
$\mathcal{L}=\mathcal{L}_{\mathcal{R}}+\alpha\mathcal{L}_{\mathcal{C}},$ (10)
where the hyperparameters $\alpha$ are used to balance the relative contributions of these two kinds of losses. We empirically set the $\alpha$ to 0.1 based on our experiment results. The ablation study of hyperparameter $\alpha$ presents in Section 4.7.2.
input :
1 $\mathcal{D}$, $\tau$     dataset and the transformations;
2 $\theta$, $s_{\theta}$    initial parameters for encoder, encoder;
3 $\xi$, $d_{\xi}$    initial parameters for decoder, decoder;
4 $\iota$, $\mathcal{H}$    initial parameters for aggregation and projector heads;
5 $\operatorname{optimizer}$    optimizer, updated parameters via gradient backwards;
6 $K$, $B$, $\mu$    pre-training steps, batch size and learning rate; for $k\leftarrow 1$ to $K$ do
       // sample one batches from dataset
7       $\mathcal{B}$leftarrow $\left\{\left\{x\right\}\sim\mathcal{D}\right\}_{i=1}^{B}$ for i, $x\in\mathcal{B}$ do
8
9            Apply multi-level masked volume generation for $x$ to obtain $x^{1}$, $x^{2}$ and $x^{3}$;
10             Apply Eq.(2) for $x^{1}$ to obtain $\mathcal{L}_{R}^{1}$ with $s_{\theta}$ and $d_{\xi}$;
11             Apply Eq.(2) for $x^{2}$ to obtain $\mathcal{L}_{R}^{2}$ with $s_{\theta}$ and $d_{\xi}$;
12             Apply Eq.(2) for $x^{3}$ to obtain $\mathcal{L}_{R}^{3}$ with $s_{\theta}$ and $d_{\xi}$;
13             Apply Eq.(3) to obtain $\mathcal{L}^{R}$ with $s_{\theta}$ and $d_{\xi}$;
14             Apply Eq.(6) for $x^{2}$ and $x^{1}$ to obtain $\mathcal{L}_{C}^{1,2}$ with $s_{\theta}$ and $\mathcal{H}_{\iota}$;
15             Apply Eq.(6) for $x^{3}$ and $x^{2}$ to obtain $\mathcal{L}_{C}^{2,3}$ with $s_{\theta}$ and $\mathcal{H}_{\iota}$;
16             Apply Eq.(7) to obtain $\mathcal{L}^{C}$ with $s_{\theta}$ and $\mathcal{H}_{\iota}$;
17             Apply Eq.(10) to obtain the final loss $\mathcal{L}$
18      $\delta\theta\leftarrow\frac{1}{B}\sum_{i=1}^{B}\left(\partial_{\theta}\mathcal{L}\right);$ $\delta\xi\leftarrow\frac{1}{B}\sum_{i=1}^{B}\left(\partial_{\xi}\mathcal{L}\right);$
19       $\delta\iota\leftarrow\frac{1}{B}\sum_{i=1}^{B}\left(\partial_{\iota}\mathcal{L}\right);$
20       $\theta\leftarrow\operatorname{optimizer}(\theta,\delta_{\theta},\mu);$ $\xi\leftarrow\operatorname{optimizer}(\xi,\delta_{\xi},\mu);$
21       $\iota\leftarrow\operatorname{optimizer}(\iota,\delta_{\iota},\mu);$
22
output : $s_{\theta}$ // the pre-trained backbone encoder
23
Algorithm 1 MiM: Mask in Mask Self-Supervised Pre-training framework
4 Experiments
Table 1: The details of each clinical dataset in our pre-training datasets.
Datasets Region of Interest
Pre-training
#Samples Pre-processing
1k 10k
BTCV [43] Abdomen 24 1. Change the volumetric orientation based on the given settings, e.g., RAS
TCIA Covid19 [2] Chest 722
LUNA16 [27] Chest 843 2. Clip intensity value to [-1000, 1000] and normalize to the range of [0, 1].
STOIC 2021 [54] Chest 2,000
FLARE23 Abdomen 4,000 3. Pad each dimension with 0 to meet the minimum shape of input.
LiDC [3] Chest 589
HNSCC [25] Head/Neck 1,071 4. Randomly crop the sub-volume with the specific shape $H_{1}\times W_{1}\times D_{1}$
TotalSegmentator [71]
Head/Neck/Chest/leg
Abdomen/pelvis/feet
1,203
Total 10,502
This section will commence by introducing the pre-training datasets, downstream datasets, and evaluation metrics. Subsequently, we will elaborate on the implementation details of our MiM. Lastly, we will present the experimental results of MiM in comparison to other SOTA methods on various tasks and datasets, along with an analysis of our proposed approach.
4.1 Datasets and Evaluation
Pre-training datasets. To conduct a fair comparison with previous works [16, 62, 34, 70, 78, 91, 90, 89, 94], we also carried out pre-training experiments on two public datasets, i.e., BTCV [43] and TCIA Covid19 [17], and combined them to form a new dataset named 1k. Additionally, to explore the scaling ability of our proposed method compared to previous state-of-the-art methods [16, 62], we collected eight publicly accessible 3D medical image datasets consisting of 10,502 CT scans to establish our pre-training datasets, which we named 10k. It is important to note that 10k is only used for exploring the scaling ability of our proposed method, while we mainly focus on the 1k dataset for fair comparison and analysis of our proposed method. Table 1 provides a summary of the sources of each collected dataset. Our pre-training process solely utilizes volumetric images without involving any annotations from any of these datasets.
Downstream datasets. To evaluate the effectiveness of our MiM, we conduct experiments on thirteen public datasets, i.e, BTCV [43], MM-WHS [37], Spleen [58], Flare22 [50], Amos22 [37], MSD Task03 [58], MSD Task06 [58], MSD Task07 [58], MSD Task08 [58], MSD Task10 [58], Covid-19-20 [56], BrasTS 21 [58], and CC-CCII [82]. These datasets include segmentation and classification tasks, with the first ten datasets used for organ segmentation, the eleventh dataset used for lesion segmentation, the twelfth dataset used for tumor segmentation, and the last dataset used for disease classification. Table 2 provides a summary of the details of each clinical dataset used in our downstream tasks. It’s noted that BTCV [43] is used in pre-training, the other datasets are unseen in pre-training. Furthermore, to evaluate the cross-modality generalization ability, we transferred the model pre-trained on CT scans to the MRI dataset BrasTS 21 [58]. We adopted consistent settings as previous work [16, 34, 62, 94].
Table 2: The details of each clinical datasets in our downstream datasets.
Datasets
Regions of
Interest
Modality #Samples Pre-processing
BTCV [43] Chest CT 30
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via HU range [-175, 250]
3. Random flip, rotation and intensities shifting with probabilities of 0.1, 0.1 and 0.5
MM-WHS [95] Heart CT 20
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via HU range [0, 1700]
Spleen [58] Spleen CT 41
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via [-57, 175]
3. Random flip, rotation and intensities shifting with probabilities of 0.5, 0.25 and 0.5
Amos22 [95] Chest CT 301
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via [-175, 250]
3. Random flip, rotation and intensities shifting with probabilities of 0.5, 0.5 and 0.5
Flare22 [95] Chest CT 50
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via [-175, 250]
3. Random flip, rotation and intensities shifting with probabilities of 0.2, 0.2 and 0.1
MSD Task03 [58] Liver CT 125
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via HU range [-21,189]
3. Random flip, rotation and intensities shifting with probabilities of 0.2, 0.2 and 0.1
MSD Task06 [58] Lung CT 63
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via [-1000, 1000]
3. Random flip, rotation and intensities shifting with probabilities of 0.5, 0.3 and 0.1
MSD Task07 [58] Pancreas CT 281
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via HU range [-87, 199]
3. Random flip, rotation and intensities shifting with probabilities of 0.5, 0.25 and 0.5
MSD Task08 [58]
Hepatic
Vessel
CT 303
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via HU range [0, 230]
3. Random flip, rotation and intensities shifting with probabilities of 0.5, 0.25 and 0.5
MSD Task10 [58] Colon CT 126
[…]
2. Crop the foregrounds with $96\times 96\times 96$ and normalize via HU range [-57, 175]
3. Random flip, rotation and intensities shifting with probabilities of 0.5, 0.25 and 0.5
Covid-19-20 [56]
Covid-19
lesion
CT 189
[…]
2. Crop the foregrounds with $192\times 192\times 32$ and normalize via HU range [-1000, 500]
BraTS 21 [58] Brain tumor MRI 484
[…]
2. Crop the foregrounds with $224\times 224\times 120$ and normalize via z-score across channels
CC-CCII [82]
Covid-19
diseases
CT 4178
Crop the foregrounds with $384\times 384\times 32$ and normalize via min-max normalization
Evaluation metrics. Following [16, 62, 94], we utilized the Dice Similarity Coefficient (DSC) [^1] https://docs.monai.io/en/latest/metrics.html#mean-dice, and Normalized Surface Distance (NSD)  [^2] https://docs.monai.io/en/latest/metrics.html#monai.metrics.SurfaceDiceMetric to assess the performance on segmentation tasks. We then utilized Accuracy (ACC) and the Area Under the Curve (AUC)  [^3] https://docs.monai.io/en/latest/metrics.html#area-under-the-roc-curve for accessing performance on disease classification.
4.2 Implementation Details and Evaluation metrics
Implementation details. In the pre-training process, we followed the settings of previous works [16, 62, 70] and provided the details of our MiM pre-training settings in Table 3. We used the backbone of  [22] as the encoder for efficient token processing. As previous SSL [4, 94], the prediction head and projection head are implemented with MLP layers for aligning the dimension. In the downstream process, we used Swin-UNETR [28] for fine-tuning on segmentation tasks and Swin-ViT for fine-tuning on classification tasks, as per the settings of previous works [62, 34, 70]. We initialized the encoder of the network with learned parameters in the pre-training process and fine-tuned the overall network. We used the AdamW [42] optimizer with a cosine learning rate scheduler for all experiments. For inference on these datasets, we applied a slicing window inference with overlapping to enable a fair comparison with previous works [62, 70]. All methods were pre-trained on four NVIDIA 3090Ti and then fine-tuned on a single NVIDIA 3090Ti. It is important to note that for evaluating the pure effectiveness of our proposed method, we did not use any foundation model or post-processing techniques [38, 47].
Table 3: Pre-training settings.
Pre-training settings
Steps 45K
Optimizer AdamW
Learning rate (LR) 1e-4
LR scheduler cosine annealing schedule
Warmup step 100
Momentum 0.9
Regularization weight 1e-2
Batch size 256
MiM Level-1 $H^{1}\times W^{1}\times D^{1}$ $384\times 384\times 192$
MiM Level-2 $H^{2}\times W^{2}\times D^{2}$ $96\times 96\times 96$
MiM Level-3 $H^{3}\times W^{3}\times D^{2}$ $16\times 16\times 16$
MiM Resize after crop $h\times w\times d$ $96\times 96\times 96$
MiM $\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{5}$ 1.0, 1.0, 1.0, 0.1, 0.1
MiM mask ratio $\tau$ 0.6
MiM sampling times on next level $\gamma$ 4
Comparison methods. In our experiments, we compared our proposed MiM method with both General SSL and Medical SSL methods. First, we compare the typical SSL methods MoCov3 [15] and MAE [16, 30], since they represent the two mainstream SSL paradigms, i.e., Contrastive learning and Pretext task. We also report the results of SimCLR [10] and PL [85], Jigsaw [8] according to  [16, 94]. We further evaluate the performance of SimMIM [79], HPM [70], localMIM [16] and MCMAE [22], since they are related to our advanced hybrid MAE. We also compare with Adam [35] due to related to our hierarchical design of multi-granularity. In addition, we compared with most existing state-of-the-art medical SSL methods in our experiments.
4.3 Experiments on downstream tasks
4.3.1 Comparison on the BTCV dataset
Outperform existing methods. We first conducted experiments on the BTCV [43], and the results are presented in Table 4. Among the comparision methods, SimCLR [14], MoCov3 [15], DINO [4], localMIM [67], HPM [67], MAE3D [16], GL-MAE [94] adopted UNETR [29] as networks architecture. Most other methods, including our MiM, used Swin-UNETR [28] as per the settings of previous works [62].
[…]
Method Pre-traning dataset Dice Score(%) Avg
Spl RKid LKid Gall Eso Liv Sto Aor IVC Veins Pan RAG LAG
From Scratch
UNETR [29] - 93.02 94.13 94.12 66.99 70.87 96.11 77.27 89.22 82.10 70.16 76.65 65.32 59.21 79.82
Swin-UNETR [28] - 94.06 93.54 93.80 65.51 74.60 97.09 75.94 91.80 82.36 73.63 75.19 68.00 61.11 80.53
with General SSL
SimCLR [10] 1k 92.79 93.04 91.41 49.65 50.99 98.49 77.92 85.56 80.58 64.37 67.16 59.04 48.99 73.85
MoCov3 [15] 1k 91.96 92.85 92.42 68.25 72.77 94.91 78.82 88.21 81.59 71.15 75.76 66.48 58.81 79.54
Jigsaw [8] 1k 94.62 93.41 93.55 75.63 73.21 95.71 80.80 89.41 84.78 71.02 79.57 65.68 60.22 81.35
[…]
SimMIM [79] 1k 95.51 93.61 93.49 67.91 73.50 96.46 81.15 89.78 84.86 72.45 75.70 66.89 64.46 81.41
[…]
with 3D Medical SSL
ROT [61] 1k 91.75 93.13 91.62 65.09 76.55 94.21 86.16 89.74 83.08 71.13 81.55 67.90 63.72 81.20
Rubik++ [92] 1k 96.21 90.41 89.33 75.22 72.64 97.44 79.25 89.65 83.76 74.74 78.35 67.14 61.97 81.38
ModelGen [91] 1k 91.99 93.52 91.81 65.11 76.14 95.98 86.88 89.29 83.59 71.79 81.62 67.97 63.18 81.45
[…]
SwinMM [70] 1k 94.33 94.18 94.16 72.97 74.75 96.37 83.23 89.56 82.91 70.65 75.52 69.17 62.90 81.81
[…]
GL-MAE [94] 1k 94.54 94.39 94.37 73.19 74.93 96.51 83.49 89.74 83.11 70.80 75.71 69.39 63.12 82.33
[…]
MAE3D [16] 1k 95.81 94.38 94.48 69.96 76.85 96.69 80.44 90.33 84.33 73.65 80.11 68.65 64.44 82.40
MiM 1k 96.05 94.58 94.53 75.73 77.36 97.03 83.23 90.37 87.64 74.97 82.62 71.02 68.80 84.46
[…]
MiM 10k 96.37 94.83 94.75 81.02 80.08 97.12 85.30 90.36 87.66 75.99 84.41 71.94 69.64 85.41
*Note: Spl: spleen, RKid: right kidney, LKid: left kidney, Gall: gallbladder, Eso: esophagus, Liv: liver, Sto: stomach, Aor: aorta, IVC: inferior vena cava, Veins: portal and splenic veins, Pan: pancreas, AG: left and right adrenal glands.
Fig. 5: Visualization of segmentation results on BTCV validation dataset [43]. We compared MiM with Unimiss [78], PCRLv2 [89], GVSL [34], Swin-UNETR [62] and MAE3D [16].
Remark. As shown in Table 4, we observed that General SSL methods performed worse than Medical SSL methods. Specifically, SimCLR [14] and MoCov3 [15] achieved only 73.85% and 79.54%, respectively. This is because these methods rely on a large batch size and negative samples to avoid trivial constants, which is impractical for 3D medical images. Additionally, the negative relationship between different images used in SimCLR [14, 11] and MoCo [31, 12, 15] is not suitable for 3D medical images. Jigsaw [8] and DINO [4] also achieved limited improvements. Our proposed method, MiM, outperformed MAE-based methods such as MAE3D [16], GL-MAE [94], localMIM [67], HPM [67], and MCMAE [22] by a significant margin. We concluded that General SSL methods are not very suitable for 3D medical images, and it is crucial to consider the characteristics of 3D medical images when designing SSL methods.
[…]
Method Pre-traning dataset BTCV MM-WHS MSD Spleen Amos22 Flare22 Avg
25% 50% 100% 25% 50% 100% 25% 50% 100% 25% 50% 100% 25% 50% 100%
From scratch
UNETR [29] - 58.99 75.20 79.82 78.77 85.31 85.85 89.67 92.23 94.20 73.53 81.46 84.34 75.32 76.29 76.94 80.53
Swin-UNETR [28] - 63.08 75.38 80.53 79.59 85.65 86.11 90.08 93.29 94.90 76.03 82.60 84.85 77.43 80.50 83.08 82.21
with General SSL
SimCLR [10] 1k 59.03 75.31 79.85 78.88 85.52 86.00 89.88 92.52 94.11 73.62 81.65 84.66 75.34 76.33 77.03 80.65
[…]
Jigsaw [8] 1k 59.83 75.89 81.35 80.51 86.50 89.75 93.23 93.61 94.29 77.43 86.09 87.49 78.01 82.85 86.40 83.55
[…]
with 3D Medical SSL
ROT [61] 1k 53.43 74.44 81.20 80.12 86.98 87.12 90.99 95.01 95.45 76.88 86.01 87.12 77.11 81.16 84.51 82.50
Rubik++ [92] 1k 53.00 74.74 81.38 80.88 87.18 88.02 91.91 95.22 95.73 77.43 86.91 88.20 78.43 82.85 86.49 83.22
ModelGen [91] 1k 54.18 61.26 81.45 80.33 86.26 89.69 92.01 95.22 95.73 77.49 86.20 88.29 78.17 82.37 84.02 82.18
[…]
GL-MAE [94] 1k 66.44 76.37 82.33 76.16 87.72 88.88 93.65 94.36 95.72 74.65 84.65 87.40 80.20 85.03 87.26 84.05
[…]
MiM 1k 70.71 81.22 84.66 86.52 88.99 91.04 95.99 96.42 96.98 80.55 88.94 90.96 85.55 86.29 88.96 87.59
[…]
MiM 10k 75.34 81.80 85.61 86.41 89.77 91.12 96.36 96.68 96.99 81.29 89.06 91.60 86.29 87.79 89.67 88.34
The scratch Swin-UNETR [28] only achieves 80.53% DSC. By pre-training with MiM on a 1k unlabeled dataset, we gained a 3.93% improvement with 84.46% DSC, which outperforms existing methods by a clear margin. Among the compared methods, Swin-UNETR [62] and MAE3D [16] achieved the best 82.58% and second-best 82.40% DSC, respectively. Our proposed MiM surpasses these two methods by 1.88% and 2.06% DSC, respectively, which is a clear improvement on this dataset.
It’s worth noting that scaling laws [41, 59, 20] also apply to 3D medical image pre-training. By pre-training with our larger scale unlabeled 10k dataset, we observed that Swin-UNETR and MAE3D achieved DSC scores of 83.20% and 83.52%, respectively. Our proposed MiM with 10k achieved a DSC score of 85.41%, which consistently outperformed these two methods by a large margin. These results suggest that scaling plays an important role in pre-training for 3D medical images and that our proposed MiM method is effective for pre-training on larger datasets.
Qualitative results. MiM was found to improve the completeness of segmentation results, as shown in Fig 5. The results of segmentation using MiM were better than those using Swin-UNETR [62], MAE3D [16], Unimiss [78], PCRLv2 [89] and GVSL [34] .
4.3.2 Comparison on the Unseen datasets
Promosing generalizability. We further conduct experiments on unseen datasets in pre-training, i.e., M-WHS [95], Spleen [58], Amos22 [37], and Flare22 [50]. The results of these four datasets are shown in Table 5. It can be observed that our MiM consistently outperformed all existing methods with a clear margin, which demonstrates promising generalizability to unseen datasets. Specifically, MiM outperformed existing methods by at least 1.89% DSC on average. By pre-training with a larger scale of unlabeled dataset 10k, Swin-UNETR [62] and MAE3D [16] improved by 0.92% and 0.94% DSC with 86.32% and 86.64%, respectively. Our MiM also gained improvement to 88.34% DSC and outperformed these two methods consistently.
Table 6: Experiment results on Task03 Liver, Task06 Lung, Task07 Pancreas, Task08 Hepatic Vessel and Task10 Colon on MSD dataset [58].
Method Pre-traning dataset Network Task03 Liver Task06 Lung Task07 Pancreas Task08 Hepatic Vessel Task10 Colon Avg
DSC(%) NSD(%) DSC(%) NSD(%) DSC(%) NSD(%) DSC(%) NSD(%) DSC(%) NSD(%) DSC(%) NSD(%)
From scratch
3D UNet [55] - - 94.41 93.94 48.09 42.65 75.33 90.98 56.11 77.58 40.97 54.63 62.98 71.95
UNETR [29] - - 94.27 94.00 44.26 40.74 74.12 88.79 58.45 79.34 27.39 42.55 59.70 69.08
Swin-UNETR [28] - - 94.52 94.08 51.97 48.68 75.75 88.87 57.24 77.84 44.57 56.15 64.84 73.12
with General SSL
MoCov3 [15] 1k UNETR 94.15 93.22 52.11 48.16 74.16 78.13 58.11 78.13 43.99 55.13 64.90 72.55
localMIM [67] 1k UNETR 94.72 93.69 52.93 48.75 74.74 79.77 58.86 78.77 44.52 54.35 65.55 73.10
HPM [67] 1k UNETR 94.63 93.25 53.63 48.34 74.74 78.13 59.89 78.93 42.67 55.13 65.51 72.81
SimMIM [79] 1k Swin-UNETR 94.50 92.45 56.93 52.41 74.99 79.77 60.44 78.99 42.42 54.35 66.26 73.65
MCMAE [22] 1k Swin-UNETR 95.49 94.33 57.08 57.18 75.78 78.13 60.44 78.93 45.63 55.13 67.28 74.96
with 3D Medical SSL
ModelGen [91] 1k 3D Unet 94.49 94.22 56.26 51.80 76.92 91.70 58.28 78.34 47.15 58.08 66.68 74.83
PCRLv2 [89] 1k 3D Unet 94.61 95.78 62.57 59.67 77.02 92.67 53.63 76.25 39.78 53.50 65.52 75.57
GL-MAE [94] 1k UNETR 94.65 93.48 53.28 49.51 72.96 87.90 60.26 78.65 45.52 59.45 65.53 74.20
MAE3D [16] 1k UNETR 95.53 96.62 58.30 52.32 77.31 91.92 60.44 78.80 45.40 56.00 66.76 75.33
Rubik++ [92] 1k Swin-UNETR 94.49 94.22 56.43 52.84 76.14 90.88 60.42 78.71 40.49 50.78 65.63 73.69
GVSL [34] 1k Swin-UNETR 95.62 95.78 59.44 54.59 77.82 92.61 57.74 77.93 50.55 61.42 68.23 76.76
Swin-UNETR [62] 1k Swin-UNETR 95.55 96.61 60.23 57.22 76.93 90.88 60.44 79.08 43.21 55.43 68.98 75.16
MiM 1k UNETR 96.01 97.03 63.06 60.11 78.39 92.71 60.71 79.44 51.02 62.57 69.79 78.37
MiM 1k Swin-UNETR 96.14 97.43 63.52 60.92 78.86 92.80 60.85 79.55 51.02 63.03 70.07 78.75
[…]
Fig. 6: Visualization of segmentation results for four tasks on the MSD dataset [58]. We compared MiM with MAE3D [16], and Swin-UNETR [62]. The Dice Score (%) is provided for each segmented volume.
Qualitative results. To gain a better understanding of the improvements achieved by MiM, we present the visualization of segmentation results for four tasks on the MSD dataset [58]. We compared our MiM method with MAE3D [16] and Swin-UNETR [62]. It can be observed in Fig. 6 that MiM outperforms MAE3D [16] and Swin-UNETR [62], by effectively addressing the issues of under and over-segmentation outliers observed in the segmentation results obtained by these methods. Specifically, Swin-UNETR [62] exhibits under-segmentation due to the absence of certain parts of the labels such as the liver and hepatic vessel, leading to a lower DSC. On the other hand, in the case of the pancreas, over-segmentation is observed with MAE3D [16] and Swin-UNETR [62], where surrounding tissues are included with the label. In contrast, our proposed MiM method effectively delineates the boundary of the pancreas.
4.4 Comparison on the Covid-19-20 dataset
Generalization capacity on Lesion Segmentation task. To evaluate the generalizability of our MiM on the lesion segmentation task, we conducted experiments on the Covid-19-20 [82] dataset. Lesion segmentation involves identifying specific abnormalities, unlike organ segmentation that focuses on delineating anatomical structures [58, 82]. showed that our proposed MiM method consistently exhibited performance improvements compared to other SSL methods by at least 1.52% and 1.73% in terms of DSC and NSD, respectively, which is a clear improvement on this dataset. By pre-training with a larger scale of unlabeled dataset 10k, MiM further improved to 71.11% and 72.42% in DSC and NSD, respectively. These results demonstrate that MiM is effective for lesion segmentation tasks.
[…]
Method Pre-traning dataset Network Lesion segmentation
DSC(%) NSD(%)
From scratch
SegResNet [51] - - 61.38 61.73
3D UNet [55] - - 60.35 61.27
UNETR [29] - - 58.80 57.22
Swin-UNETR [28] - - 61.89 61.92
with General SSL
MoCov3 [15] 1k UNETR 63.15 64.32
localMIM [67] 1k UNETR 66.62 67.11
simMIM [79] 1k Swin-UNETR 67.22 68.19
[…]
with Medical SSL
Adam [35] 1k SegResNet 67.31 67.56
ModelGen [91] 1k 3D UNet 68.25 69.63
PCRLv1 [90] 1k 3D UNet 67.25 68.14
PCRLv2 [89] 1k 3D UNet 68.18 69.18
GL-MAE [94] 1k UNETR 67.11 68.94
[…]
Rubik++ [92] 1k Swin-UNETR 64.60 65.00
[…]
MiM 1k Swin-UNETR 70.14 71.02
[…]
4.5 Comparison on CC-CCII dataset
Generalization capacity on Disease Classification task. To evaluate the generalizability of our MiM on classification task, we fine-tuned it on the CC-CCII [82] dataset and compared its performance with state-of-the-art general and medical SSL methods. Since existing SSL methods didn’t conduct experiments on this dataset, we reproduced the related methods and reported the results. As shown in Table 8, it can be observed that MiM achieved the best performance in terms of ACC and AUC, surpassing all other methods with 93.63% and 99.39%. These findings demonstrate that the learned representation by MiM can be well transferable to classification problems and can be used effectively for medical image classification tasks. With a larger scale of pre-training dataset 10k, our MiM further improved to 94.12% and 99.52% in ACC and AUC, respectively, which shows the scalability of our proposed method when transferring across tasks.
Table 8: Experiment results on CC-CCII dataset [82].
Method Pre-traning dataset Network Disease classification
ACC(%) AUC
From Scratch
ResNet [32] - - 85.94 87.61
ViT [18] - - 85.92 96.26
Swin-ViT [48] - - 87.15 97.32
with General SSL
SimCLR [10] 1k ResNet 87.12 96.21
MoCov2 [12] 1k ResNet 87.01 95.49
localMIM [67] 1k ViT 88.15 97.58
HPM [66] 1k ViT 88.26 97.65
Jigsaw [8] 1k Swin-ViT 87.18 96.89
SimMIM [79] 1k Swin-ViT 89.62 98.16
MCMAE [22] 1k Swin-ViT 90.26 97.12
with Medical SSL
PCRLv1 [90] 1k ResNet 88.84 97.69
PCRLv2 [89] 1k ResNet 89.35 98.05
GL-MAE [94] 1k ViT 88.00 96.97
MAE3D [16] 1k ViT 91.30 98.13
Rubik++ [92] 1k Swin-ViT 89.93 98.55
SwinMM [70] 1k Swin-ViT 89.99 98.77
Swin-UNETR [62] 1k Swin-ViT 91.81 98.97
MiM 1k Swin-ViT 93.63 99.39
[…]
4.6 Comparison on the BraTS 21 dataset
Generalization capacity on MRI modality dataset. To evaluate the generalizability of our MiM on MRI datasets, we fine-tuned it on the BraST 21 [58] MRI tumor segmentation dataset and compared its performance with state-of-the-art general and medical SSL methods. It can be observed in Table 9 that both SSL methods can improve the performance of the model in segmenting tumors in BrasTS 21 [58] dataset. This is because CT and MRI are often used for the same task but for different purposes, and thus share similar anatomical structures. Therefore, the knowledge learned by SSL methods from unlabeled CT datasets can be transferred to MRI datasets [49, 62]. Our MiM outperformed all other methods by at least 1.34% with 79.28% DSC. By pre-training with a larger unlabeled dataset 10k, our MiM further improved to 79.92%, which shows the scalability of our proposed method when transferring across modalities.
[…]
Method Pre-traning dataset Network Dice Score(%)
TC WT ET AVG
From Scratch
UNETR [29] - - 81.62 87.81 57.34 75.58
Swin-UNETR [28] - - 81.28 88.67 57.73 75.89
with General SSL
SimCLR [10] 1k UNETR 83.13 89.44 58.42 76.99
MoCov3 [15] 1k UNETR 82.60 88.89 57.69 76.39
[…]
Jigsaw [8] 1k Swin-UNETR 81.62 89.45 59.10 76.72
[…]
with 3D Medical SSL
GL-MAE [94] 1k UNETR 82.80 90.65 60.46 77.97
[…]
PCRLv1 [90] 1k Swin-UNETR 81.96 88.83 57.58 76.12
PCRLv2 [89] 1k Swin-UNETR 82.13 90.06 57.70 76.63
Rubik++ [92] 1k Swin-UNETR 84.32 90.23 58.01 77.51
Swin-UNETR [62] 1k Swin-UNETR 82.51 89.08 58.15 76.58
SwinMM [70] 1k Swin-UNETR 83.48 90.47 58.72 77.56
MiM 1k Swin-UNETR 84.99 91.92 60.94 79.28
[…]
Label efficiency. The scarcity of labeled 3D medical images can pose a significant challenge, requiring models to perform well despite this constraint [62]. To evaluate the effectiveness of MiM under this constraint, we compared it with existing methods by reducing the number of manual labels at label ratios of 25% and 50%. As shown in Table 5. MiM outperformed all other methods by a significant margin when generalized to both BTCV [43] and unseen datasets. Specifically, MiM with 50% label achieved comparable performance with the scratch Swin-UNETR with 100% label with a clear margin.
4.6.1 Comparison on the MSD datasets
Generalizability to Organ Segmentation. To evaluate the generalizability, we conducted experiments on the five CT-based tasks on the MSD dataset [58], i.e., Task 03 Liver, Task06 Lung, Task07 Pancreas, Task08 Hepatic Vessel and Task10 Colon. Since existing methods didn’t conduct experiments with the same pre-training dataset, we reimplemented these methods for fair comparison. It can be observed in Table 6 that MiM achieves the best average DSC (70.07%) and NSD(78.75%) in all tasks. Since the scrtach Swin-UNETR [28] performs better UNETR [29] in terms of avg DSC (64.84% vs 62.98%) and NSD (73.12% vs 69.08). We further pre-trained UNETR [29] based on MiM for a fair comparison. It can be observed that with MiM pre-training, Swin-UNETR [28] gained an average improvement 5.23% and 5.63% in terms of DSC and NSD, respectively. With UNETR [29] as the network, we observed an average improvement of 10.09% and 9.29% in terms of DSC and NSD, respectively. Moreover, by pre-training with a larger scale of unlabeled dataset 10k, MiM further improved to 70.76% and 79.67% in DSC and NSD, respectively. These results demonstrate that MiM is effective for organ segmentation tasks.
4.7 Analysis of our proposed method
We first conducted ablation studies on loss functions and settings of MiM. Then, we conducted a hyper-parameter analysis for the loss terms. Lastly, we provided reconstruction results. All the models were pre-trained on 1k, which were then evaluated on BTCV [43] and MM-WHS [95].
4.7.1 Ablation study
Loss functions. To investigate the impact of hierarchical design by applying multi-level reconstruction and cross-level alignment, we conducted a comprehensive ablation study of the loss terms. Table 10 presents the results of the ablation studies on three validation datasets. By utilizing multi-level reconstruction, the model significantly improved upon the previous MiM method, as the model was able to capture multi-granularity visual cues of the 3D medical images and learn representation hierarchically, thereby learning broader information. The cross-level alignment for coarse $x^{1}$ and $x^{2}$ (4th row) achieved better performance than that of fine-cropped $x^{3}$ 3D medical images (5th row). Combining all losses could result in best performance improvements.
Table 10: Evaluation of the loss terms $\mathcal{L}_{R}$ and $\mathcal{L}_{C}$. We report the performance on BTCV [43], MM-WHS [95], and Spleen [58] dataset.
Loss BTCV MM-WHS
$\mathcal{L}^{\mathcal{R}}_{1}$ $\mathcal{L}^{\mathcal{R}}_{2}$ $\mathcal{L}^{\mathcal{R}}_{3}$ $\mathcal{L}^{\mathcal{C}}_{12}$ $\mathcal{L}^{\mathcal{C}}_{23}$ DSC(%) NSD(%) DSC NSD(%)
81.49 75.92 89.42 73.06
83.10
79.51
90.44
74.53
84.25
81.29
90.74
75.50
[…]
84.29
81.45
90.89
75.97
84.43
81.56
90.89
76.27
84.66
82.11
91.04
76.65
Numbers of levels. We limit the number of levels $L$ to three. As illustrated in Fig. 7, we varied the number of levels $L$ and observed that the optimal performance was achieved when $L=3$. This choice stems from the fact that the resolution of the volume at Level-4 is considerably smaller in scale compared to the original volume ($1\times 1\times 1$), facilitating easier reconstruction but resulting in suboptimal performance [30, 79]. Moreover, the proximity of positive patches to negative patches within a confined region may hinder effective discrimination, particularly concerning hard negative patches [40]. Thus, we set the number of levels $L=3$ as the default choice.
Fig. 7: Evaluation of the number of level $L$ for our multi-level volume generation. We report the results on BTCV [43] and MM-WHS [95].
Reconstruction targets for $\mathcal{L}_{\mathcal{R}}$. In Fig. 3, we employ the mask patches from the coarser as the finer level volume, which is the default choice in MiM. In Table 11, we observed that adopting mask patches from the coarser level as the finer level volume results in better performance than the other choices.
Table 11: Evaluation of the process of finer level volume. We report the DSC on BTCV [43] and MM-WHS [95].
Finer level volume BTCV MM-WHS
DSC(%) NSD(%) DSC(%) NSD(%)
Visible patches 82.18 76.22 89.33 73.65
Maked patches 84.66 82.11 91.04 76.65
Negative pairs for $\mathcal{L}_{\mathcal{C}}$. In Eq. 6, we employ infoNCE loss[10] as the default choice. This loss maximizes the similarity between the cross-level images and pushes away the negative samples. Another loss function is BYOL cosine loss [24]. The primary distinction is whether to use negative samples. As demonstrated in Table 12, negative samples aid in learning better representations [12], and thus, we include them in our default choices.
Table 12: Evaluation of $\mathcal{L}_{C}$ whether using negative samples. We report the DSC on BTCV [43] and MM-WHS [95].
Loss function $L_{C}$ BTCV MM-WHS
DSC(%) NSD(%) DSC(%) NSD(%)
BYOL-style cosine loss [24] 83.68 80.12 90.53 74.66
Contrastive loss [10] 84.66 82.11 91.04 76.65
Efficiency of MiM. Table 13 presents a comparative analysis of computation costs i.e. flops and times during the pre-training period across various methodologies. Our evaluation highlights the computational efficiency of MAE-based approaches when coupled with UNETR, as they leverage only the visible tokens from masked 3D medical images. In contrast, alternative MAE-based methods naively cooprating with hybrid backbones such as Swin-UNETR, necessitate processing all tokens. Our proposed method extends the hybrid backbone architecture [22] to 3D medical images in the pre-training stage, achieving significantly superior performance while maintaining computational efficiency.
Table 13: Evaluation of model complexity during pre-training and subsequent performance on downstream tasks. We report flops and pre-training duration for pre-training. We report Dice Score(%) on BTCV [43] and MM-WHS [95] dataset for fine-tuning.
Method Networks Tokens flops(G) Pre-training duration BTCV MM-WHS
All Partial
MAE3D UNETR 12.3 9 82.20 90.03
HPM UNETR 34.1 9.9 82.04 89.86
localMIM UNETR 26.6 10.3 81.38 89.99
Adam Swin-UNETR 43.7 27.0 80.45 88.82
SimMIM Swin-UNETR 70.8 26.6 82.24 90.22
Swin-UNETR Swin-UNETR 394.84 130.0 82.58 90.32
MiM Swin-UNETR 45.8 14.6 84.66 91.04
4.7.2 Hyper-parameter analysis
Weights of $\mathcal{L}_{\mathcal{R}}$ and $\mathcal{L}_{\mathcal{C}}$. In Eq. 10, the loss function of MiM consists of two parts, namely multi-level reconstruction and cross-level alignment. To determine the suitable class weight for these loss terms, we varied the value of $\alpha$ in Fig. 8. When we varied the value of $\alpha$ from a small value of $1e^{-2}$ to larger value $1e^{-1}$ during the pre-training period, it strengthened the impact of cross-level alignment between finer views and larger views, thereby enforcing the representation of local and global alignment. As a result, the DSC of the MiM fine-tuned on the BTCV validation dataset improved by $0.42\%$. However, further scaling up the weight of cross-level alignment did not bring any additional benefits. This may be because the magnitude of the cross-level alignment function is much larger than the magnitude of the reconstruction loss. Thus, scaling up the weight of cross-level alignment leads the model to ignore the reconstruction process, which is also crucial for MiM method.
Fig. 8: Evaluation of the loss weight $\alpha$ for $\mathcal{L_{C}}$. We report the results on BTCV dataset [43].
Weight of different levels of $\mathcal{L}_{\mathcal{R}}$. To determine the appropriate weights of each level for MiM, we applied different weights to determine the priority of the level of masked 3D medical images during the pre-training, as shown in Fig. 1. Specifically, we explored focusing on reconstructing the coarse level of masked images first, then the finer level (2nd row), and reconstructing the finer level first, then the coarse level of masked images (3rd row) by dynamically updating the weight with a cosine scheduler, as presented in Table 14. The results demonstrated that learning from finer to coarser masked images would help learn a more discriminative representation. Furthermore, we explored simultaneously learning the multi-granularity visual clues (4th row), which outperformed the other settings and set as default choice.
Table 14: Evaluation of the model trained with different learning paradiagm by varying the weight of each level of $\mathcal{L}_{\mathcal{R}}$, repectively. We reprot Disce Score(%) on BTCV [43] and MM-WHS [95] dataset.
Learning process Weight Datasets
Level-1 Level-2 Level-3 BTCV MM-WHS
w/o Coarse & Fine 1$\longrightarrow$1 0$\longrightarrow$0 0$\longrightarrow$0 81.49 89.42
[…]
Simultaneously 1$\longrightarrow$1 1$\longrightarrow$1 84.66 91.04
4.7.3 Reconstruction results
We provided reconstruction results of MiM on BTCV [43] in Fig. 9, where the first and second rows represent the volumes and masked 3D medical images, respectively, and the reconstruction results in the last row demonstrate excellent performance.
Fig. 9: Reconstruction results of the proposed method MiM on the BTCV validation dataset [43].
5 Discussion
5.1 Limitations
While our MiM framework has shown promising results, our MiM does have limitations. In our experiments, we only scaled up the MiM to a 10k pre-training dataset, which may not be sufficient to fully explore its potential. Previous works in Natural Language Processing [41] and Computer Vision [41, 68] have demonstrated that scaling up the pre-training dataset to billions of images can lead to the foundation model for their fields [41], but collecting enough high-quality 3D medical images remains challenging. Therefore, further exploration is needed to investigate the performance of MiM with a vast scale of pre-training datasets, such as 100k.
Moreover, while the hierarchical design is effective, we adopted a simple design in the experiments to consider computation costs. For the generation of multi-level volumes, we employed a random sampling strategy to select the masked patches as next-level volumes, which achieved good performance. Since not all patches are equally important [45], Wang et al. [66] proposed a hard negative mining strategy to sample the patches for reconstruction. Future exploration can design priory sampling strategy i.e., selecting the most informative patches for finer level reconstruction, to further improve the performance of MiM. Finally, as our MiM framework mainly mines the hierarchical structure for intra-volume relationships, it can be extended to explore its capacity in mining inter-volume relationships, which may further improve its performance.
5.2 Future directions and challenges
The foundation model has the potential to advance healthcare [33], and SSL is a promising direction for building the foundation model in 3D medical images [33, 1]. However, there are current limitations that need to be addressed. To build upon the 3D medical image foundation model with SSL, several promising emerging directions can be explored. (1) Building up a large-scale pre-training datasets, e.g., more than 100k volumes. (2) Exploring more downstream tasks, e.g., 3D medical image registration [86], and tumor synthesis [36]. (3) Exploring cooperation with other modalities for multi-modal pre-training, e.g., language [73, 1, 44]. Actually, (1) is the cornerstone for the foundation model in 3D medical images, and (2) can further comprehensively evaluate the performance of SSL for 3D medical image analysis. Since existing SSL methods often pre-train with a limited scale of data and evaluate on limited downstream tasks, which cannot lead to a healthcare foundation model for 3D medical images [73]. Inspired by the success of  Achiam et al. [1], Radford et al. [53], multi-modal pre-training for 3D medical image analysis also deserves more attention, which can learn robust representation by complementing the information from different modalities.
6 Conclusion
In this paper, we introduce the Mask in Mask (MiM) pre-training framework, which significantly advances 3D medical image analysis. MiM improves MAE by facilitating the learning of hierarchical visual tokens across various scales. This results in a robust representation for transfer learning in downstream application-specific tasks. By incorporating hierarchical designs, i.e., multi-level reconstruction and cross-level alignment, MiM efficiently encodes multi-granularity visual cues of structure and details into the representation. To facilitate the fair and comprehensive comparison of existing methods, we collected ten public datasets and curated two scales of pre-training datasets, i.e., 1k and 10k. Consistent with existing works, we pre-trained our MiM on a 1k dataset, and extensive experiments on thirteen datasets demonstrate MiM’s effectiveness in achieving superior performance over other methods, including segmentation and classification tasks. We also transferred the learned representation of MiM to MRI-based tasks to evaluate the cross-modality generalization ability. Additionally, we conducted an ablation study to validate the importance of each component of MiM. We conducted a hyper-parameter analysis to investigate the impact of different hyper-parameters on the performance of MiM. The results reveal that the hierarchical design of the MiM framework is crucial for achieving superior performance for 3D medical images. We further explore to scale up the pre-training dataset to 10k. The results show that the performance of MiM can be further improved by scaling up the pre-training dataset. This finding emphasizes the importance of large-scale pre-training for building the foundation model in 3D medical images. The study hopes to serve as a strong baseline for future research in 3D medical image analysis and draw more attention from the community to the importance of hierarchical design and large-scale pre-training towards the healthcare foundation model.
7 Acknowledgments
This work was supported by the Hong Kong Innovation and Technology Fund (Project No. MHP/002/22), Shenzhen Science and Technology Innovation Committee Fund (Project No. SGDX20210823103201011), and the Project of Hetao Shenzhen-Hong Kong Science and Technology Innovation Cooperation Zone (HZQB- KCZYB-2020083).